OpenStudy (anonymous):

find the direction of w, given w=5i-3j

OpenStudy (tkhunny):

How shall we report it? Maybe an angle from the positive x-axis?

OpenStudy (anonymous):

im not quite sure what that means, or how to go about it

OpenStudy (tkhunny):

What if it were a point on a cartesian coordinate system. Which way do you have to look to find (5,-3) when you are standing at the Origin?

OpenStudy (anonymous):

to the left in quadrant II

OpenStudy (tkhunny):

?? No, it's in Quadrant IV.

OpenStudy (anonymous):

OHH right right my bad wow

OpenStudy (anonymous):

Perhaps the question is asking for the unit vector

OpenStudy (tkhunny):

Maybe: Or \(\theta\) in the 4th Quadrant, given \(tan(\theta) = -3/5\)

OpenStudy (anonymous):

what if it is for vectors, because thats what this chapter is all about

OpenStudy (tkhunny):

Well, how are you reporting vectors? You get only Magnitude and Direction. What are the units for "Direction"?

OpenStudy (anonymous):

It's probably the unit vector. You can way w is a vector of length |w| in the w/|w| direction

OpenStudy (tkhunny):

As good a choice as any.